In the Michelson experiment to measure c, light from a lamp source is reflected from one of eight mirrored faces of an octagonal wheel toward a mirror 35 km away. The light is reflected back to the wheel, where it is reflected into the observer's eye, as shown. What minimum angular speed should the wheel have so that the observer sees the light from the lamp? What is the next highest speed for which this is possible?

A: 3364 s^-1 ; 6729 s^-1
Answer #1

From Experiment, w angular frequency Steady image for 6 7 I speed of light distance travelled by the light between reflection

G= 3x108 m/s D D = 2x35 - 70 km - 70ooom, o = 216 DN 27 (3x108/ 70000 x 8 w = 3366 rads Next highest speed =W col = 20 W!.